# Discrete Approximate Circle Bundles

### v2

@@ 2.3 Classification Of Circle Bundles
The projection map $\widetilde{\exp}:\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})\to\text{Homeo}_{+}(\mathbb{S}^{1})$ is defined by $\widetilde{\exp}(\tilde{f})=\text{exp}\circ\tilde{f}\circ\text{exp}^{-1}$, where $\exp:\mathbb{R}\to\mathbb{S}^{1}$ is given by $\exp(\theta)=e^{2\pi i\theta}$.
Notice that sequence (2) splits, and the section $r:\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ induced by complex conjugation gives rise to a semidirect decomposition $\text{Homeo}(\mathbb{S}^{1})\cong\text{Homeo}_{+}(\mathbb{S}^{1})\rtimes_{r}\mathbb{Z}_{2}$. In more detail, we obtain a right $\mathbb{Z}_{2}$-action on $\text{Homeo}_{+}(\mathbb{S}^{1})$ via the assignment $\pm 1\longmapsto\text{Adj}_{r(\pm 1)}$ which induces a group multiplication law $(\Lambda_{1},\omega_{1})\cdot(\Lambda_{2},\omega_{2})=(\Lambda_{1}\text{Adj}_{r(\omega_{1})}(\Lambda_{2}),\omega_{1}\omega_{2})$. The map $\varphi:\text{Homeo}_{+}(\mathbb{S}^{1})\rtimes_{r}\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ given by $\varphi(\Lambda,\omega)=\Lambda r(\omega)$ is then a group isomorphism.
Thus, for any open cover $\mathcal{U}$ of a paracompact space $B$, $\varphi$ induces an isomorphism on cohomology, where the cocycle condition is given explicitly by $(\Lambda_{jk},\omega_{jk})\cdot(\Lambda_{kl},\omega_{kl})=(\Lambda_{jl},\omega_{jl})$, or, equivalently,
$\Lambda_{jk}\text{Adj}_{r(\omega_{jk})}\Lambda_{kl}=\Lambda_{jl}\quad\quad\omega_{jk}\omega_{kl}=\omega_{jl}\\$
Consequently, for any $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$, there is a natural bijection between cocycles $\Omega\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})$ satisfying $\deg_{*}(\Omega)=\omega$ and cochains $\Lambda\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1})})$ satisfying $\Lambda_{jk}\text{Adj}_{r(\omega_{jk})}\Lambda_{kl}=\Lambda_{jl}$ for all $(jkl)\in\mathcal{N}(\mathcal{U})$. Alternatively, we can interpret $\omega$ as a local system and $\Lambda$ as a Cech cocycle in $\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})$:
@@
Suppose $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ for some open cover $\mathcal{U}$ of a paracompact space $B$. The following sequence is short exact:
@@
Moreover, the sheaf $\mathcal{C}_{\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1}),\omega}$ is soft, so the induced connecting homomorphism $\Delta_{\omega}:\check{H}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$ is an isomorphism.
@@
At the level of stalks, sequence (6) reduces to exact sequence (3), hence it is also exact as a sequence of sheaves of abelian groups. By Proposition 2.2, we therefore obtain a long exact sequence in Cech cohomology, and in particular, a connecting homomorphism $\Delta_{\omega}:\check{H}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$. Moreover, $\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})$ is contractible, so $\mathcal{C}_{\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1}),\omega}$ is soft, so Proposition 2.2 implies $\Delta_{\omega}$ is an isomorphism by exactness.
The proposition above immediately leads to the following classification result:
@@
For any open cover $\mathcal{U}$ of a paracompact space $B$, there is a bijective correspondence
via the assignment $\Omega]\longmapsto\left(w_{1}(\Omega),\tilde{e}(\Omega)\right)$ given by $w_{1}(\Omega)=[\deg_{*}(\Omega)]$ and $\tilde{e}(\Omega)=\Delta_{\omega}([\Omega])$.
We therefore have a complete classification of circle bundles over paracompact spaces in terms of a pair of discrete invariants. As we will later show, these two invariants are closely connected to the first Stiefel-Whitney class and Euler class for real plane bundles, so we will refer to $w_{1}(\Omega)$ and $\tilde{e}(\Omega)$ respectively as the Stiefel-Whitney class (or orientation class) and the twisted Euler class of the circle bundle represented by $\Omega$.
Remark 2.12.
The notation in Proposition 2.11 is slightly abusive, since each cocycle $\omega$ within a cohomology class determines a different local system $\mathbb{Z}_{\omega}$. On the other hand, by Proposition 2.8, the local systems determined by any two cohomologous cocycles are isomorphic. In particular, if $\omega$ is a coboundary, then $\mathbb{Z}_{\omega}\cong\underline{\mathbb{Z}}$.
Observe that when $B$ is a locally finite 1-dimensional CW complex (e.g., a graph), the Proposition 2.11 implies that circle bundles over $B$ are completely determined by their Stiefel-Whitney classes. In the case when $B$ is a compact connected surface, the classification is also especially simple and explicit:
Proposition 2.13.
Let $B$ be a closed connected 2-manifold, and let $\omega\in H^{1}(B;\mathbb{Z}_{2})$ determine a double cover $p_{\omega}:B_{\omega}\to B$. Then
@@
Twisted Poincare duality asserts that for any compact connected $n$-manifold $M$ and locally constant sheaf $\mathcal{L}$ of Abelian groups on $M$ (with fiber $G$), one has $H^{*}(M;\mathcal{L})\cong H_{n-*}(M;\mathcal{L}\otimes\mathcal{O})$, where $\mathcal{O}$ denotes the orientation sheaf of $M$. Moreover, $H_{0}(M;\mathcal{L})$ is isomorphic to the coinvariants of the monodromy action $\omega_{\mathcal{L}}:\pi_{1}(M)\to\text{Aut}(G)$ induced by $\mathcal{L}$, i.e.,
@@
Thus, for any compact connected surface $B$ and orientation character $\omega\in H^{1}(B;\mathbb{Z}_{2})$, one has
@@
Since $\mathbb{Z}_{\omega}\otimes\mathcal{O}\cong\underline{\mathbb{Z}}$ if and only if $\omega+w_{1}(B)=0$ (or, equivalently, $\omega=w_{1}(B)^{-1}=w_{1}(B)$), the proposition follows.
In light of the proposition above, for any compact connected surface $B$ and $\omega\in Z^{1}(B;\mathbb{Z}_{2})$, one can define a twisted fundamental class $B]_{\omega}$ which generates $\check{H}_{0}(B;\mathbb{Z}_{\omega})$ – this choice is unique when $\omega\neq w_{1}(B)$ and unique up to sign when $\omega=w_{1}(B)$. The pairing $\langle\tilde{e},[B]_{\omega}\rangle$ then gives an explicit bijection between $\check{H}^{2}(B;\mathbb{Z}_{\omega})$ and its image in $\mathbb{Z}$. We call the integer $\langle\tilde{e}(\Omega),[B]_{\omega}\rangle$ the twisted Euler number of the bundle classified by $\Omega$.
@@ 2.4 Reduction Of Structure Group
The structure group $\text{Homeo}(\mathbb{S}^{1})$ is rather unwieldly – in particular, it is disconnected, non-abelian and infinite-dimensional, making it challenging to work with from a computational perspective. In this section, we identify reductions of this structure group to simpler ones.
@@
    The cohomology class $\Omega_{\xi}\in\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})$ lifts to a class in $\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1})})$,
@@
    The Stiefel-Whitney class $w_{1}(\Omega_{\xi})\in\check{H}^{1}(B;\underline{\mathbb{Z}_{2}})$ vanishes.
@@
The inclusion $\iota:\mathbb{S}^{1}\to\text{Homeo}_{+}(\mathbb{S}^{1})$ is a homotopy equivalence.
@@
Consider the map $\tau:\text{Homeo}_{+}(\mathbb{S}^{1})\to\mathbb{S}^{1}$ defined by $\tau(f)=f(1)$. Observe that $\tau$ is a continuous surjection, and the fiber of $\tau$ over each $z\in\mathbb{S}^{1}$ contains the orientation-preserving homeomorphisms which map $1$ to $z$. One can easily check that each $\tau^{-1}(z)$ is contractible, so $\tau$ is a homotopy inverse for $\iota$.
@@
More generally, Proposition 2.15 implies that the inclusion of $\mathbb{S}^{1}\rtimes\mathbb{Z}_{2}=\text{Iso}(\mathbb{S}^{1})$ into $\text{Homeo}(\mathbb{S}^{1})$ is a homotopy equivalence. It follows that over paracompact spaces, any circle bundle, no matter how wild, admits a family of trivializations with $\text{Iso}(\mathbb{S}^{1})$-valued transition maps.
While the inclusion $\iota:\mathbb{S}^{1}\to\text{Homeo}_{+}(\mathbb{S}^{1})$ induces a one-to-one correspondence between oriented circle bundles and principal $\mathbb{S}^{1}$-bundles, we note that the retract $\tau:\text{Homeo}_{+}(\mathbb{S}^{1})\to\mathbb{S}^{1}$ defined in the proof of Proposition 2.15 is not a group homomorphism, hence does not induce a cochain map $\check{C}^{1}(B;\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1})})\to\check{C}^{1}(B;\mathcal{C}_{\mathbb{S}^{1}})$ which descends to an isomorphism on cohomology. In particular, the following proposition shows that no retract can induce such a map:
Proposition 2.16.
There does not exist a retract $\tau:\text{Homeo}_{+}(\mathbb{S}^{1})\to\mathbb{S}^{1}$ which is also a group homomorphism.
Proof.
One can show $\text{Homeo}_{+}(\mathbb{S}^{1})$ is a perfect group in the sense that every element is expressible as a product of commutators (see [Ghy01] Proposition 5.11). On the other hand, if $\tau:\text{Homeo}_{+}(\mathbb{S}^{1})\to\mathbb{S}^{1}$ is a group homomorphism, we must have $\tau([f,g])=[\tau(f),\tau(g)]$ for all $g,h\in\text{Homeo}_{+}(\mathbb{S}^{1})$. Since $\mathbb{S}^{1}$ is Abelian and $\text{Homeo}_{+}(\mathbb{S}^{1})$ is perfect, it follows immediately that $\tau(f)=1$ for all $f$. We conclude that $\tau$ cannot be a retract onto $\mathbb{S}^{1}$.
Thus, given an arbitrary family of trivializations for a circle bundle, it may be difficult to find explicit homotopies so that the new transitions maps take values in $\text{Iso}(\mathbb{S}^{1})$. On the other hand, if we define the supremum distance on $\text{Homeo}(\mathbb{S}^{1})$ by
$d_{\infty}(f,g)=\sup_{z\in\mathbb{S}^{1}}d_{\mathbb{S}^{1}}(f(z),g(z))\\$ (7)
where $d_{\mathbb{S}^{1}}$ denotes the geodesic distance metric on $\mathbb{S}^{1}$, the proposition below states that well-behaved homeomorphisms of $\mathbb{S}^{1}$ are close to isometries. Since the proof requires some technical digressions, we defer it to the appendix.
Proposition 2.17.
If $f\in\text{Homeo}(\mathbb{S}^{1})$ is $L$-Lipschitz, then $d_{\infty}(f,\text{Iso}(\mathbb{S}^{1}))\leq 2\pi(\frac{L-1}{L})$.
Note in particular that $d_{\infty}(f,\text{Iso}(\mathbb{S}^{1}))\to 0$ as $L\to 1$.
From a computational perspective, the structure group $\text{Iso}(\mathbb{S}^{1})\cong\mathbb{S}^{1}\rtimes\mathbb{Z}_{2}$ is vastly preferable to $\text{Homeo}(\mathbb{S}^{1})$. In particular, $\text{Iso}(\mathbb{S}^{1})$ is a 1-dimensional Lie group (with two connected components), and it has a natural identification with the matrix group $O(2)$ which gives rise to the following commutative diagram of split extensions:
@@
where $\exp:\mathbb{R}\to SO(2)$ is given by $\exp(\theta)=\left(\begin{array}[]{cc}\cos(2\pi\theta)&-\sin(2\pi\theta)\\ \sin(2\pi\theta)&\cos(2\pi\theta)\\ \end{array}\right)$. Passing to Cech cohomology, it is clear that the Stiefel-Whitney and twisted Euler classes we defined to classify circle bundles correspond precisely to their classical counterparts for vector bundles. However, it is useful to know how these classes arise and can be explicitly computed from arbitrary $\text{Homeo}(\mathbb{S}^{1})$-valued transition data.
2.5 The Universal Circle Bundle And Dimensionality Reduction
A classifying map can be used to model the total space of a fiber bundle as an embedded submanifold of a more familiar space. Specifically, if $G$ is a topological group and $\mu_{G}=(EG,p,BG)$ is a universal $G$-bundle, then for any principal $G$-bundle $\xi$ over a paracompact space $B$, we can construct a classifying map $f:B\to BG$ such that such that $f^{*}\mu_{G}=\xi$. This map is covered by a bundle map $F:E(\xi)\to EG$ which embeds the total space of $\xi$ into $EG$. More generally, if $\xi_{F}$ is the associated bundle of $\xi$ with fiber $F$, then $\xi_{F}$ is isomorphic to the pullback of the associated bundle $EG\times_{G}F\to BG$ along the same classifying map.

### v3

@@ 2.3 Classification Of Circle Bundles
[…]
Notice that sequence (2) splits, and the section $s:\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ induced by complex conjugation, gives rise to a semidirect decomposition $\text{Homeo}(\mathbb{S}^{1})\cong\text{Homeo}_{+}(\mathbb{S}^{1})\rtimes_{s}\mathbb{Z}_{2}$. Explicitly, we obtain a right $\mathbb{Z}_{2}$-action on $\text{Homeo}_{+}(\mathbb{S}^{1})$ via the assignment $\pm 1\longmapsto\text{Adj}_{s(\pm 1)}$ defined by $\text{Adj}_{s(\pm 1)}(\Lambda)=s(\pm 1)\Lambda s(\pm 1)^{-1}$. This induces a group multiplication law $(\Lambda_{1},\omega_{1})\cdot(\Lambda_{2},\omega_{2})=(\Lambda_{1}\text{Adj}_{s(\omega_{1})}(\Lambda_{2}),\omega_{1}\omega_{2})$. The map $\varphi:\text{Homeo}_{+}(\mathbb{S}^{1})\rtimes_{s}\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ given by $\varphi(\Lambda,\omega)=\Lambda s(\omega)$ is then a group isomorphism.
Thus, for any open cover $\mathcal{U}$ of a paracompact space $B$, $\varphi$ induces an isomorphism on cohomology, where the cocycle condition $(\Lambda_{jk},\omega_{jk})\cdot(\Lambda_{kl},\omega_{kl})=(\Lambda_{jl},\omega_{jl})$ is given explicitly by
$\Lambda_{jk}\text{Adj}_{r(\omega_{jk})}(\Lambda_{kl})=\Lambda_{jl}\quad\quad\omega_{jk}\omega_{kl}=\omega_{jl}\\$
Consequently, for any $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$, there is a natural bijection between cocycles $\Omega\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})$ satisfying $\deg_{*}(\Omega)=\omega$ and cochains $\Lambda\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1})})$ satisfying $\Lambda_{jk}\text{Adj}_{r(\omega_{jk})}(\Lambda_{kl})=\Lambda_{jl}$ for all $(jkl)\in\mathcal{N}(\mathcal{U})$. Alternatively, we can interpret $\omega$ as a local system and $\Lambda$ as a Čech cocycle in $\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})$:
@@
Suppose $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ for some open cover $\mathcal{U}$ of a paracompact space $B$. The following sequence of sheaves is short exact:
@@
Moreover, the sheaf $\mathcal{C}_{\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1}),\omega}$ is soft, so the induced connecting homomorphism
$\Delta_{\omega}:\check{H}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$
is an isomorphism.
@@
At the level of stalks, sequence (6) reduces to exact sequence (3), hence it is also exact as a sequence of sheaves of abelian groups. By Proposition 2.3, we therefore obtain a long exact sequence in Čech cohomology, and in particular, a connecting homomorphism $\Delta_{\omega}:\check{H}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$. Moreover, $\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})$ is contractible, so $\mathcal{C}_{\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1}),\omega}$ is soft, and Proposition 2.3 implies $\Delta_{\omega}$ is an isomorphism by exactness.
The proposition above immediately leads to the main classification result of this section:
Theorem 2.12.
For any open cover $\mathcal{U}$ of a paracompact space $B$, there is a natural bijective correspondence
via the assignment $\Omega]\longmapsto\left(\mathsf{w}_{1}(\Omega),\tilde{e}(\Omega)\right)$ given by $\mathsf{w}_{1}(\Omega)=[\deg_{*}(\Omega)]$ and $\tilde{e}(\Omega)=\Delta_{\omega}([\Omega])$.
We therefore have a complete classification of circle bundles over paracompact spaces in terms of a pair of discrete invariants. As we will later show, these two invariants are closely connected to the first Stiefel-Whitney class and Euler class for real plane bundles, so we will refer to $\mathsf{w}_{1}(\Omega)$ and $\tilde{e}(\Omega)$ respectively as the Stiefel-Whitney class (or orientation class) and the twisted Euler class of the circle bundle represented by $\Omega$.
Remark 2.13.
The notation in Theorem 2.12 is slightly abusive, since each cocycle $\omega$ within a cohomology class determines a different local system $\mathbb{Z}_{\omega}$. On the other hand, by Proposition 2.9, the local systems determined by any two cohomologous cocycles are isomorphic. In particular, if $\omega$ is a coboundary, then $\mathbb{Z}_{\omega}\cong\underline{\mathbb{Z}}$.
Observe that when $B$ is a locally finite 1-dimensional CW complex (e.g., a graph), then Theorem 2.12 implies that circle bundles over $B$ are completely classified by the first Stiefel-Whitney class. In the case when $B$ is a compact connected surface, the classification is also especially simple and explicit:
@@
Let $B$ be a closed connected 2-manifold, and let $\omega]\in H^{1}(B;\mathbb{Z}_{2})$. Then
@@
Twisted Poincare duality 2.10 asserts that for any compact connected $n$-manifold $M$ and locally constant sheaf $\mathcal{L}$ of Abelian groups on $M$ (with fiber $G$), one has $H^{*}(M;\mathcal{L})\cong H_{n-*}(M;\mathcal{L}\otimes\mathcal{O})$, where $\mathcal{O}$ denotes the orientation sheaf of $M$. Moreover, $H_{0}(M;\mathcal{L})$ is isomorphic to the coinvariants of the monodromy action $\omega_{\mathcal{L}}:\pi_{1}(M)\to\text{Aut}(G)$ induced by $\mathcal{L}$, i.e.,
@@
Thus, for any compact connected surface $B$ and orientation character $\omega]\in H^{1}(B;\mathbb{Z}_{2})$, one has
@@
Since $\mathbb{Z}_{\omega}\otimes\mathcal{O}\cong\underline{\mathbb{Z}}$ if and only if $\omega]+\mathsf{w}_{1}(B)=0$ (or, equivalently, $\omega]=\mathsf{w}_{1}(B)^{-1}=\mathsf{w}_{1}(B)$), the proposition follows.
In light of the proposition above, for any compact connected surface $B$ and $\omega]\in H^{1}(B;\mathbb{Z}_{2})$, one can define a twisted fundamental class $B]_{\omega}$ which generates $H_{2}(B;\mathbb{Z}_{\omega})$ – this choice is unique when $\omega]\neq\mathsf{w}_{1}(B)$ and unique up to sign when $\omega]=\mathsf{w}_{1}(B)$. The pairing $\langle\tilde{e},[B]_{\omega}\rangle$ then gives an explicit bijection between $\check{H}^{2}(B;\mathbb{Z}_{\omega})$ and its image in $\mathbb{Z}$. We call the integer $\langle\tilde{e}(\Omega),[B]_{\omega}\rangle$ the twisted Euler number of the bundle classified by $\Omega$.
@@ 2.4 Reduction Of Structure Group
The group $\text{Homeo}(\mathbb{S}^{1})$ is disconnected, non-abelian and infinite-dimensional, making it challenging to work with from a computational perspective. In this section, we identify reductions of this structure group to simpler ones.
@@
    The cohomology class $\Omega_{\xi}]\in\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})$ lifts to a class in $\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1})})$,
@@
    The orientation class $\mathsf{w}_{1}([\Omega_{\xi}])\in\check{H}^{1}(B;\underline{\mathbb{Z}_{2}})$ vanishes.
@@
The inclusion $\iota:\mathbb{S}^{1}\to\text{Homeo}_{+}(\mathbb{S}^{1})$ given by multiplication in $\mathbb{S}^{1}$ is a homotopy equivalence.
@@
Consider the map $\tau:\text{Homeo}_{+}(\mathbb{S}^{1})\to\mathbb{S}^{1}$ defined by $\tau(f)=f(1)$. Observe that $\tau$ is a continuous surjection, and the fiber of $\tau$ over each $z\in\mathbb{S}^{1}$ is comprised of the orientation-preserving homeomorphisms which map $1$ to $z$. One can check that each $\tau^{-1}(z)$ is contractible, and a theorem of Smale [28] then implies that $\tau$ is a homotopy inverse for $\iota$.
@@
More generally, Proposition 2.16 implies that the inclusion of $\mathbb{S}^{1}\rtimes\mathbb{Z}_{2}\cong\text{Iso}(\mathbb{S}^{1})$ into $\text{Homeo}(\mathbb{S}^{1})$ is a homotopy equivalence (where $\text{Iso}(\mathbb{S}^{1})$ denotes isometries of $\mathbb{S}^{1}$ with respect to geodesic distance). It follows that over paracompact spaces, any circle bundle, no matter how wild, admits a family of trivializations with $\text{Iso}(\mathbb{S}^{1})$-valued transition maps.
From a computational perspective, the structure group $\text{Iso}(\mathbb{S}^{1})\cong\mathbb{S}^{1}\rtimes\mathbb{Z}_{2}$ is vastly preferable to $\text{Homeo}(\mathbb{S}^{1})$. In particular, $\text{Iso}(\mathbb{S}^{1})$ is a 1-dimensional Lie group (with two connected components), and it has a natural identification with the group $O(2)$ of $2\times 2$ orthogonal matrices, which gives rise to the following commutative diagram of split extensions:
@@
[…]
2.5 The Tautological Circle Bundle Over $\text{Gr}(2)$ And Dimensionality Reduction
A classifying map can be used to identify the total space of a fiber bundle with a subspace of a more familiar object. Specifically, if $G$ is a topological group and $\mu_{G}=(EG,p,BG)$ is a universal $G$-bundle, then for any principal $G$-bundle $\xi=(E,\pi,B)$ over a paracompact space $B$, we can construct a classifying map $f:B\to BG$ with $f^{*}\mu_{G}\cong\xi$. This map is covered by a bundle map $\tilde{f}:E\to EG$ in the sense that $p\circ\tilde{f}=f\circ\pi$, and serves to map the total space of $\xi$ into $EG$. More generally, if $\xi^{\prime}=(E\times_{G}F,\pi^{\prime},B)$ is the associated bundle of $\xi$ with fiber a left $G$-space $F$, then $\xi^{\prime}$ is isomorphic to the pullback of the associated bundle $EG\times_{G}F\to BG$ along the same classifying map $f$. In this sense, we think of $EG\times_{G}F\to BG$ as “universal”, though the total space is generally not contractible.